# Generalize across Homophily and Heterophily: Hybrid Spectral Graph Pre-Training and Prompt Tuning

### v2

@@ 5.1 Experimental Setup
Datasets. We conduct experiments on real-world datasets with various homophily levels. Among them, Cora (McCallum et al. 2000), Citeseer (Sen et al. 2008), and Pubmed (Kipf and Welling 2016) are homophilic graphs, Cornell, Texas, Wisconsin (Pei et al. 2020), Chameleon, Squirrel (Rozemberczki, Allen, and Sarkar 2021), Amazon-ratings (Ratings) and Roman-empire (Empire) (Platonov et al. 2023) are considered as heterophilc graphs. The details of datasets are in Appendix B.3 and Table 4.
Baselines. The baselines fall into three categories: (i) Traditional GNN models: GCN (Kipf and Welling 2016), GAT (Veličković et al. 2017), BernNet (He et al. 2021), ChebNetII (He, Wei, and Wen 2022) and TFE-GNN (Duan et al. 2024); (ii) Graph “pre-training and fine-tuning” models: DGI (Veličković et al. 2018), GraphCL (You et al. 2020), SimGRACE (Xia et al. 2022) and PolyGCL (Chen, Lei, and Wei 2024); (iii) Graph “pre-training and prompt tuning” models: GPPT (Sun et al. 2022), GPrompt (Liu et al. 2023), GPF-plus (Fang et al. 2024), All-in-One (Sun et al. 2023), and ProNoG (Yu et al. 2024). Among them, TFE-GNN, PolyGCL, and ProNoG are designed to handle heterophily. For DGI, GraphCL, SimGRACE, GPPT, GPrompt, GPF-plus, and All-in-One, we use GCN (Kipf and Welling 2016) as the backbone model, while other backbones are evaluated in Appendix E.5. Details are in Appendix B.1.
Implementation Details. We focus on two scenarios: transductive and inductive settings. In the former, the pre-training graph is the same as the downstream graphs, while in the latter, it’s different. Following prior graph prompt work (Sun et al. 2023; Yu et al. 2024), we define our task as 5-shot node classification. We set $C$ in the hybrid spectral filter backbone to 2, resulting in 3 different filters. The prompt node number in each prompt graph is 10. Due to label imbalance, we use the F1 score as the metric, with accuracy results in the Appendix E.4. More details are in Appendix B.2. Besides the main text, we also conduct parameter sensitivity evaluation and runtime analysis in Appendix E.2 and E.3.
@@ 5.2 Performance Comparison
Results of transductive learning are in Table 1, inductive in Table 2. More results are in Appendix E.4 and E.5.
@@
Inductive Performance We assess the inductive learning performance across graphs with different homophily levels. We examine both in-domain (where pre-training and downstream tasks are in the same domain) and cross-domain settings. Four dataset pairs are randomly selected while more are in Appendix E.4 and E.5. The results show our model excels in inductive learning, despite the greater challenge of knowledge transfer due to the disparity between pre-trained and downstream graphs. It acquires rich spectral knowledge during pre-training and aligns the downstream graph’s spectral distribution with pre-trained knowledge during prompt tuning, enhancing knowledge transfer.
@@ 5.3 Ablation Study
We investigate the key modules’ significance in pre-training and prompt tuning. In pre-training, we introduce HS-GPPT (low-pass) with a single GCN (Kipf and Welling 2016) as GNN backbone. In prompt tuning, we develop three simplified versions: HS-GPPT (single prompt), HS-GPPT (w/o prompt), and HS-GPPT (w/o prompt norm). HS-GPPT (single prompt) uses one prompt graph for all filters. HS-GPPT (w/o prompt) omits prompt graphs, and HS-GPPT (w/o prompt norm) skips prompt graph normalization. Partial results are in Table 3, with more in Appendix E.6.
@@
Table 3: The ablation study of different variants.
@@ 6 Conclusion
In this paper, we present HS-GPPT, a novel graph “pre-training and prompt tuning” model designed to handle graphs with varying homophily levels. Existing methods, limited by the homophily assumption, struggle to generalize across graphs with diverse spectral distributions. We address this by using a hybrid spectral filter backbone and local-global contrastive learning to capture a wide range of spectral information. In prompt tuning, we introduce prompt graphs to align with spectral pre-trained knowledge for better transfer. Experiments show HS-GPPT’s superior performance over baselines across different homophily levels in both transductive and inductive settings.
@@ Appendix A Additional Proofs
A.1 Proof of Theorem 3.2
Proof. Here we denote the number of edges $|\mathcal{E}|$, therefore the number of homophilic edges is $h|\mathcal{E}|$ while the number of heterophilic edges is $(1-h)|\mathcal{E}|$. Then
@@
In spectral graph theory (Spielman 2007; Sandryhaila and Moura 2014; Tang et al. 2022), with the unnormalized graph Laplacian matrix $\mathbf{L}=\mathbf{D}-\mathbf{A}$, $\mathbf{x}^{T}\mathbf{L}\mathbf{x}=\mathbf{x}^{T}\mathbf{D}\mathbf{x}-\mathbf{x}^{T}\mathbf{A}\mathbf{x}=\sum_{{(u,v)}\in\mathcal{E}}{(x_{u}-x_{v})}^{2}$, then we have:
@@
$=|\mathcal{E}|\cdot\Big{\{}h\cdot\mathcal{E}(d_{intra})+(1-h)\cdot\mathbb{E}(d_{inter})\Big{\}}$
@@
Generally, the inter-class distance is greater than the intra-class distance, thus we have $\mathbb{E}(d_{inter})>\mathbb{E}(d_{intra})$. Therefore $S_{high}$ is monotonically decreasing with $h$.
@@
Proof. Previous work (Chen, Lei, and Wei 2024) proves that for normalized $\hat{\mathbf{x}}$ where $\sum_{i=1}^{N}\hat{x}_{i}^{2}=N$, the Spectral Regression Loss satisfies
@@
where $c$ denotes the assumed upper bound of $g(\lambda_{i})$ and acts as a constant term in this context. Thus, the upper bound of $\mathcal{L}(\mathcal{G})$ is:
@@
Discarding constant terms, maximizing $\mathcal{L}_{UB}(\mathcal{G})$ reduces to
@@
Set $w_{i}=\hat{x}_{i}^{2}/N$, so that $w_{i}\geq 0$ and $\sum_{i}w_{i}=1$. Writing $g_{i}=g(\lambda_{i})$, the objective becomes
$\sum_{i=1}^{N}\hat{x}_{i}^{2}\,g_{i}=N\sum_{i=1}^{N}w_{i}\,g_{i}.$ (21)
Hence it suffices to maximize the linear functional $\sum_{i}w_{i}g_{i}$ over the probability simplex. Suppose, for the sake of contradiction, that in an optimal weight vector $\mathbf{w}^{*}$ there exist indices $i\neq j$ with
@@
Choose $0<\varepsilon\leq w^{*}_{j}$ and define a perturbed weight vector $\mathbf{w}^{\prime}$ by
@@
Then $\mathbf{w}^{\prime}$ remains feasible and
@@
Since $g_{i}-g_{j}>0$,
$\sum_{k=1}^{N}w^{\prime}_{k}\,g_{k}>\sum_{k=1}^{N}w^{*}_{k}\,g_{k},$ (25)
which contradicts the maximality of $\mathbf{w}^{*}$. Hence, in any maximizer one must have
@@
Recalling $w_{i}=\hat{x}_{i}^{2}/N$, this shows that larger values of $g(\lambda_{i})$ must be paired with larger $\hat{x}_{i}^{2}$, which completes the proof.
@@
Here $\mathbf{D}$ is related to $\mathbf{A}$. Therefore, when Inequality 29 is satisfied, the following three scenarios exist:
@@
  • $\mathbf{X}_{1}\neq\mathbf{X}_{2},\mathbf{A}_{1}\neq\mathbf{A}_{2}$. In such case, both the adjacent matrix and node features of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are different. There, we can transform $\mathcal{G}_{1}$ into $\mathcal{G}_{2}$ through a combination of transformation operations (e.g., “changing node features”, “adding or deleting edges”, and “adding or removing isolated sub-graphs”) (Fang et al. 2024).
@@
Then we introduce Lemma A.2 from (Sun et al. 2023):
@@
Following prior work (Fang et al. 2024; Sun et al. 2023), the effect of introducing prompt nodes can be equivalently viewed as a transformation on the features of the original nodes. Since our analysis ultimately focuses on the performance on these original nodes, we therefore maintain a fixed node number in our theoretical framework, modeling the prompt’s effect as a change from the original signal $\hat{\mathbf{x}}$ to a new signal $\hat{\tilde{\mathbf{x}}}$. Therefore, the SRL of the two graphs are:
@@
In Theorem 4.1, we prove that the prompt graph can simulate any spectral distribution. Thereore, there exists a prompt graph $\mathcal{G}_{p}$ such that
@@
Combine Equation 33, 34 and Inequality 35, we have:
@@ B.1 Baselines Descriptions
  • GCN (Kipf and Welling 2016): This method achieves neighborhood information aggregation by spectral graph convolutions.
  • GAT (Veličković et al. 2017): This method proposes an attention mechanism to aggregate neighborhood information.
  • BernNet (He et al. 2021): This spectral method overcomes the limitations of oversimplified or ill-posed filters. It estimates any filter over a graph’s normalized Laplacian spectrum via an order-K Bernstein polynomial approximation and designs spectral properties by setting Bernstein basis coefficients.
  • ChebNetII (He, Wei, and Wen 2022): This spectral method is based on Chebyshev interpolation, which enhances the original Chebyshev polynomial approximation while reducing the Runge phenomenon.
  • TFE-GNN (Duan et al. 2024): This spectral method generalizes both homophily and heterophily through ensembles of multiple spectral filters.
@@
  • DGI (Veličković et al. 2018): DGI functions as a self-supervised pre-training method designed for homogeneous graphs. It is based on the maximization of mutual information (MI), with the aim of increasing the estimated MI between locally augmented instances and their global equivalents.
  • GraphCL (You et al. 2020): GraphCL utilizes a range of graph augmentations for self-supervised learning, exploiting the inherent structural patterns of graphs. The main objective is to enhance the consistency between different augmentations during graph pre-training.
  • SimGRACE (Xia et al. 2022): SimGRACE overcomes the limitations of existing GCL methods related to data augmentations. It utilizes the original graph and a perturbed GNN model as encoders to generate contrastive views without the need for data augmentations. Additionally, it employs an adversarial training scheme to boost robustness, thereby attaining competitive performance along with high flexibility and efficiency.
  • PolyGCL (Chen, Lei, and Wei 2024): PolyGCL is spectral-based and solves the limitations of current methods when dealing with heterophilic graphs. It employs polynomial filters to produce spectral views for contrastive learning between low-pass and high-pass views, incorporating high-pass information. This enables PolyGCL to show superiority on graphs with varying homophily.
@@
  • GPPT (Sun et al. 2022): GPPT utilizes the link prediction task to pre-train a GNN model and reformulates the downstream node classification task into the link prediction task. However, it has been proven that the link prediction task only preserves the low-frequency similarity information (Liu et al. 2022; Yu et al. 2024).
  • GPrompt (Liu et al. 2023): GraphPrompt employs subgraph similarity prediction to pre-train the GNN model and reformulates the downstream node and graph classification into the subgraph similarity prediction template. However, the subgraph similarity prediction task is similar to the link prediction task, which only captures low-pass information.
  • GPF-plus (Fang et al. 2024): GPF proposes a universal prompt-based tuning method for pre-trained GNN models, regardless of the pre-training strategy. It operates within the input graph’s feature space and theoretically attains an equivalent effect to any prompting function, thereby eliminating the need for an explicit illustration of prompting functions for each strategy. Here, we utilize the more flexible and robust version, GPF-plus.
  • All-in-One (Sun et al. 2023): All-in-One reformulates all the tasks into sub-graph classification and adopts prompt graphs based on GPF-plus (Fang et al. 2024) to adapt universally to any pre-training strategy.
  • ProNoG (Yu et al. 2024): ProNoG addresses the limitations of existing graph prompt methods for non-homophilic graphs. ProNoG first analyzes existing graph pre-training methods to provide theoretical insights on pre-training task choices, and then proposes a conditional network to characterize node-specific non-homophilic patterns in downstream tasks.

### v3

@@ 5.1 Experimental Setup
Datasets. We conduct experiments on real-world datasets with various homophily levels. Among them, Cora McCallum et al. (2000), Citeseer Sen et al. (2008), and Pubmed Kipf and Welling (2016) are homophilic graphs, Cornell, Texas, Wisconsin Pei et al. (2020), Chameleon, Squirrel Rozemberczki et al. (2021), Amazon-ratings (Ratings) and Roman-empire (Empire) Platonov et al. (2023) are considered as heterophilic graphs. The details of datasets are in Appendix B.3 and Table 4.
Baselines. The baselines fall into three categories: (i) Traditional GNN models: GCN Kipf and Welling (2016), GAT Veličković et al. (2017), BernNet He et al. (2021), ChebNetII He et al. (2022) and TFE-GNN Duan et al. (2024); (ii) Graph “pre-training and fine-tuning” models: DGI Veličković et al. (2018), GraphCL You et al. (2020), SimGRACE Xia et al. (2022) and PolyGCL Chen et al. (2024a); (iii) Graph “pre-training and prompt tuning” models: GPPT Sun et al. (2022), GPrompt Liu et al. (2023), GPF-plus Fang et al. (2024), All-in-One Sun et al. (2023), and ProNoG Yu et al. (2024). Among them, TFE-GNN, PolyGCL, and ProNoG are designed to handle heterophily. For DGI, GraphCL, SimGRACE, GPPT, GPrompt, GPF-plus, and All-in-One, we use GCN Kipf and Welling (2016) as the backbone model, while other backbones are evaluated in Appendix E.4. Details are in Appendix B.1.
Implementation Details. We focus on two scenarios: transductive and inductive settings. In the former, the pre-training graph is the same as the downstream graphs, while in the latter, it’s different. Following prior graph prompt work Sun et al. (2023); Yu et al. (2024), we define our task as 5-shot node classification. We set $C$ in the hybrid spectral filter backbone to 2, resulting in 3 different filters. The prompt node number in each prompt graph is 10. Due to label imbalance, we use the F1 score as the metric, with accuracy results in the Appendix E.3. More details are in Appendix B.2. Besides the main text, we also conduct parameter sensitivity evaluation and runtime analysis in Appendix E.1 and E.2.
@@ 5.2 Performance Comparison
Results of transductive learning are in Table 1, inductive in Table 2. More results are in Appendix E.3 and E.4.
@@
Inductive Performance We assess the inductive learning performance across graphs with different homophily levels. We examine both in-domain (where pre-training and downstream graphs are in the same domains) and cross-domain settings (where pre-training and downstream graphs are in different domains). Four dataset pairs are randomly selected, while more are in Appendix E.3 and E.4. The results show our model excels in inductive learning, despite the greater challenge of knowledge transfer due to the disparity between pre-trained and downstream graphs. It acquires rich spectral knowledge during pre-training and aligns the downstream graph’s spectral distribution with pre-trained knowledge during prompt tuning, enhancing knowledge transfer.
@@ 5.3 Ablation Study
We investigate the key modules’ significance in pre-training and prompt tuning. In pre-training, we introduce HS-GPPT (low-pass) with a single GCN Kipf and Welling (2016) as GNN backbone. In prompt tuning, we develop three simplified versions: HS-GPPT (single prompt), HS-GPPT (w/o prompt), and HS-GPPT (w/o prompt norm). HS-GPPT (single prompt) uses one prompt graph for all filters. HS-GPPT (w/o prompt) omits prompt graphs, and HS-GPPT (w/o prompt norm) skips prompt graph normalization. Partial results are in Table 3, with more in Appendix E.5.
@@
Table 3: The ablation study of different variants. More results are in Appendix E.5.
@@ 6 Conclusion
In this paper, we identify two fundamental limitations in existing graph prompt tuning methods: the knowledge bottleneck caused by single-filter backbones, and the utilization bottleneck arising from the mismatch between pre-trained spectral knowledge and downstream spectra. To overcome this, we propose HS-GPPT, establishing an abundantt spectral basis via hybrid pre-training. Crucially, we treat prompt tuning as a spectral alignment process, employing learnable prompts to align with spectral pre-trained knowledge for fully utilization and better transfer. Experiments show HS-GPPT’s superior performance over baselines across different homophily levels in both transductive and inductive settings.
@@ Appendix A Additional Proofs
A.1 Proof of Proposition 3.2
Proof. Given a graph $\mathcal{G}$ with node label $\mathbf{Y}$ and homophily level $h$, we denote the expectation of normalized intra-class and inter-class distances Ju et al. [2025] as:
$\small\mathbb{E}(d_{intra})=\mathbb{E}\left[\frac{(x_{u}-x_{v})^{2}}{\mathbf{x}^{T}\mathbf{x}}\right]_{e_{uv}\in\mathcal{E},y_{u}=y_{v}}$ (12)
$\mathbb{E}(d_{inter})=\mathbb{E}\left[\frac{(x_{u}-x_{v})^{2}}{\mathbf{x}^{T}\mathbf{x}}\right]_{e_{uv}\in\mathcal{E},y_{u}\neq y_{v}},$ (13)
where $\mathbb{E}(d_{intra})$ and $\mathbb{E}(d_{inter})$ reflect the node feature distribution. In general, the distance between nodes of different classes is larger than that within the same class, implying $\mathbb{E}(d_{inter})>\mathbb{E}(d_{intra})$.
Let $|\mathcal{E}|$ denote the total number of edges. Consequently, the number of homophilic edges is $h|\mathcal{E}|$ and the number of heterophilic edges is $(1-h)|\mathcal{E}|$. By expanding the expectations, we have:
@@
In spectral graph theory Spielman [2007]; Tang et al. [2022], using the regular graph Laplacian $\mathbf{L}=\mathbf{D}-\mathbf{A}$, the quadratic form is $\mathbf{x}^{T}\mathbf{L}\mathbf{x}=\mathbf{x}^{T}\mathbf{D}\mathbf{x}-\mathbf{x}^{T}\mathbf{A}\mathbf{x}=\sum_{{(u,v)}\in\mathcal{E}}{(x_{u}-x_{v})}^{2}$. The high-frequency energy $S_{high}$ is derived as:
@@
$=|\mathcal{E}|\cdot\Big\{h\cdot\mathbb{E}(d_{intra})+(1-h)\cdot\mathbb{E}(d_{inter})\Big\}$
@@
With $\mathcal{C}_{base}={|\mathcal{E}|}\mathbb{E}(d_{inter})$ and $\mathcal{C}_{gap}={|\mathcal{E}|}[\mathbb{E}(d_{inter})-\mathbb{E}(d_{intra})]$, we have:
$S_{high}=\mathcal{C}_{base}-\mathcal{C}_{gap}\cdot h.$ (18)
Since $\mathbb{E}(d_{inter})>\mathbb{E}(d_{intra})$ and $|\mathcal{E}|>0$, we have $\mathcal{C}_{gap}>0$. Therefore, $S_{high}$ is monotonically decreasing with $h$.
@@
Proof. First, we formally introduce the optimization objective. Following Chen et al. [2022], we focus on binary node classification, as multi-class problems can be decomposed into multiple binary tasks. Let $\mathbf{U}\in\mathbb{R}^{N\times N}$ be the eigenvector matrix of the graph Laplacian. Let $\mathbf{Y}\in\mathbb{R}^{N\times 2}$ be the label matrix, with columns $\mathbf{y}_{0},\mathbf{y}_{1}$ as class indicators. and define the label difference vector $\Delta\mathbf{y}=\mathbf{y}_{0}-\mathbf{y}_{1}$. The spectral label is denoted as $\hat{\mathbf{y}}=\mathbf{U}^{T}\Delta\mathbf{y}=(\hat{y}_{1},\cdots,\hat{y}_{N})$, and the spectral signal of node features is $\hat{\mathbf{x}}=\mathbf{U}^{T}\mathbf{x}=(\hat{x}_{1},\cdots,\hat{x}_{N})$.
We employ the Spectral Regression Loss (SRL) Lei et al. [2022] to evaluate the quality of the graph filter $g$. The loss is defined as:
$\mathcal{L}(\mathcal{G})=\sum_{i=1}^{N}\left(\frac{\hat{y}_{i}}{\sqrt{N}}-\frac{g(\lambda_{i})\hat{x}_{i}}{\sqrt{\sum_{j=1}^{N}\big({g(\lambda_{j})}^{2}\hat{x}_{j}^{2}\big)}}\right)^{2}.$ (19)
A smaller SRL indicates better alignment between the filtered signal and the ground truth labels, implying better downstream performance.
Based on the derivation in previous work Chen et al. [2024a], for a normalized signal where $\sum_{i=1}^{N}\hat{x}_{i}^{2}=N$, the SRL is bounded by:
@@
where $c$ is a constant representing the upper bound of $g(\lambda_{i})$. Thus, the upper bound of $\mathcal{L}(\mathcal{G})$ is:
@@
To minimize the loss upper bound $\mathcal{L}_{UB}(\mathcal{G})$, we must maximize the subtraction term. Discarding constants, the optimization problem becomes:
@@
Let $w_{i}=\hat{x}_{i}^{2}/N$. Since $\sum_{i}w_{i}=1$ and $_{i}\geq 0, 𝐰\mathbf{w}$ lies on the probability simplex. Let $g_{i}=g(\lambda_{i})$. The objective simplifies to maximizing the linear functional $J=\sum_{i}w_{i}g_{i}$.
Suppose, for the sake of contradiction, that in an optimal weight distribution $\mathbf{w}^{*}$, the signal energy is not aligned with the filter response. specifically, there exist indices $i,j$ such that:
@@
Choose a perturbation $0<\varepsilon\leq w^{*}_{j}$ and define a new weight vector $\mathbf{w}^{\prime}$:
@@
The vector $\mathbf{w}^{\prime}$ remains feasible. Comparing the objectives:
@@
Since $g_{i}>g_{j}$ and $\varepsilon>0$, we have $\varepsilon(g_{i}-g_{j})>0$, which implies:
$\sum_{k=1}^{N}w^{\prime}_{k}\,g_{k}>\sum_{k=1}^{N}w^{*}_{k}\,g_{k}.$ (26)
This contradicts the assumption that $\mathbf{w}^{*}$ is optimal. Therefore, to maximize the objective (and minimize SRL), we must have:
@@
Recalling $w_{i}\propto\hat{x}_{i}^{2}$, this proves that larger filter responses $g(\lambda_{i})$ must be paired with larger signal energy $\hat{x}_{i}^{2}$ (i.e., positive correlation) to minimize the loss.
@@
Here $\mathbf{D}$ is related to $\mathbf{A}$. Therefore, when Inequality 30 is satisfied, the following three scenarios exist:
@@
  • $\mathbf{X}_{1}\neq\mathbf{X}_{2},\mathbf{A}_{1}\neq\mathbf{A}_{2}$. In such case, both the adjacent matrix and node features of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are different. There, we can transform $\mathcal{G}_{1}$ into $\mathcal{G}_{2}$ through a combination of transformation operations (e.g., “changing node features”, “adding or deleting edges”, and “adding or removing isolated sub-graphs”) Fang et al. [2024].
@@
Then we introduce Lemma A.2 from Sun et al. [2023]:
@@
Following prior work Fang et al. [2024]; Sun et al. [2023], the effect of introducing prompt nodes can be equivalently viewed as a transformation on the features of the original nodes. Since our analysis ultimately focuses on the performance on these original nodes, we therefore maintain a fixed node number in our theoretical framework, modeling the prompt’s effect as a change from the original signal $\hat{\mathbf{x}}$ to a new signal $\hat{\tilde{\mathbf{x}}}$. Therefore, the SRL of the two graphs are:
@@
In Theorem 4.1, we prove that the prompt graph can simulate any spectral distribution. Therefore, there exists a prompt graph $\mathcal{G}_{p}$ such that
@@
Combine Equation 34, 35 and Inequality 36, we have:
@@ B.1 Baselines Descriptions
  • GCN Kipf and Welling [2016]: This method achieves neighborhood information aggregation by spectral graph convolutions.
  • GAT Veličković et al. [2017]: This method proposes an attention mechanism to aggregate neighborhood information.
  • BernNet He et al. [2021]: This spectral method overcomes the limitations of oversimplified or ill-posed filters. It estimates any filter over a graph’s normalized Laplacian spectrum via an order-K Bernstein polynomial approximation and designs spectral properties by setting Bernstein basis coefficients.
  • ChebNetII He et al. [2022]: This spectral method is based on Chebyshev interpolation, which enhances the original Chebyshev polynomial approximation while reducing the Runge phenomenon.
  • TFE-GNN Duan et al. [2024]: This spectral method generalizes both homophily and heterophily through ensembles of multiple spectral filters.
@@
  • DGI Veličković et al. [2018]: DGI functions as a self-supervised pre-training method designed for homogeneous graphs. It is based on the maximization of mutual information (MI), with the aim of increasing the estimated MI between locally augmented instances and their global equivalents.
  • GraphCL You et al. [2020]: GraphCL utilizes a range of graph augmentations for self-supervised learning, exploiting the inherent structural patterns of graphs. The main objective is to enhance the consistency between different augmentations during graph pre-training.
  • SimGRACE Xia et al. [2022]: SimGRACE overcomes the limitations of existing GCL methods related to data augmentations. It utilizes the original graph and a perturbed GNN model as encoders to generate contrastive views without the need for data augmentations. Additionally, it employs an adversarial training scheme to boost robustness, thereby attaining competitive performance along with high flexibility and efficiency.
  • PolyGCL Chen et al. [2024a]: PolyGCL is spectral-based and solves the limitations of current methods when dealing with heterophilic graphs. It employs polynomial filters to produce spectral views for contrastive learning between low-pass and high-pass views, incorporating high-pass information. This enables PolyGCL to show superiority on graphs with varying homophily.
@@
  • GPPT Sun et al. [2022]: GPPT utilizes the link prediction task to pre-train a GNN model and reformulates the downstream node classification task into the link prediction task. However, it has been proven that the link prediction task only preserves the low-frequency similarity information Liu et al. [2022]; Yu et al. [2024].
  • GPrompt Liu et al. [2023]: GraphPrompt employs subgraph similarity prediction to pre-train the GNN model and reformulates the downstream node and graph classification into the subgraph similarity prediction template. However, the subgraph similarity prediction task is similar to the link prediction task, which only captures low-pass information.
  • GPF-plus Fang et al. [2024]: GPF proposes a universal prompt-based tuning method for pre-trained GNN models, regardless of the pre-training strategy. It operates within the input graph’s feature space and theoretically attains an equivalent effect to any prompting function, thereby eliminating the need for an explicit illustration of prompting functions for each strategy. Here, we utilize the more flexible and robust version, GPF-plus.
  • All-in-One Sun et al. [2023]: All-in-One reformulates all the tasks into sub-graph classification and adopts prompt graphs based on GPF-plus Fang et al. [2024] to adapt universally to any pre-training strategy.
  • ProNoG Yu et al. [2024]: ProNoG addresses the limitations of existing graph prompt methods for non-homophilic graphs. ProNoG first analyzes existing graph pre-training methods to provide theoretical insights on pre-training task choices, and then proposes a conditional network to characterize node-specific non-homophilic patterns in downstream tasks.